OpenStudy (anonymous):

What is the 10th term of the sequence 625, 125, 25, … ?

Miracrown (miracrown):

ok 5^4 = 625 5^3 = 125 5^2 = 25 that seems to be the pattern so the next term is 5^1 = 5